if you have a 93 sweets then how would you divide them equally into your bags. each bag contain equal number of sweets ?

Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)

you will have 31 bags,with 3 sweets in each bag
31x3========93

Answer by ikleyn(53427)

.
if you have 93 sweets then how would you divide them equally into your bags.
each bag contain equal number of sweets ?
~~~~~~~~~~~~~~~~~~~~~~~~

This problem is very special. It says "distribute sweets equally into your bags", but does not specify
how many bags are there.

So, we should analyze/represent all possible cases.


The number 93 is a composite number. It has several decompositions into the products of prime numbers

    93 = 1*93 = 3*31 = 31*3 = 93*1.


According to these multiple decompositions, there are several cases.


(1)  If there is one bag, you place there all 93 sweets into this bag..

(2)  If there are 3 bags, you place 31 sweets into each bag.

(3)  If there are 31 bags, you place 3 sweets into each bag.

(4)  If there are 93 bags, you place 1 sweet into each bag.


Thus all possible cases are covered.

Solved.

This problem is to connect the conception of prime decomposition with other
interesting ideas in your young mind - for better understanding.